OpenStudy (anonymous):

in an arithmetic sequence, the first term is 6 and the fifth term is twice the fourth term. find the common difference

hartnn (hartnn):

n'th term is given by an = a+(n-1)d

hartnn (hartnn):

so put., a5= 2a4 a=6

hartnn (hartnn):

bdw, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

so could you find common difference ? 'd' or still need help?

OpenStudy (anonymous):

i think i need a bit more help sorry

hartnn (hartnn):

sure. the n'th term of Arithmetic sequence is given by \(\large a_n=a+(n-1)d\) where a is 1st term = 6 d is common difference(which u need to find) now its given to u that a5=2a4 u can get a5 as a+4d =6+4d a4 as a+3d = 6+3d and since a5=2a4 6+4d=2(6+3d) now can u find d ?

OpenStudy (anonymous):

thanks so much